# First Integrals of Geodesic Flows on Cones

### v1

@@
Abstract
In this paper we study the behavior of geodesics on cones over arbitrary $C^{3}$-smooth closed Riemannian manifolds. We show that the geodesic flow on such cones admits first integrals whose values uniquely determine almost all geodesics except for cone generatrices. This investigation is inspired by our recent results on billiards inside cones over manifolds where similar results hold true.
@@ 1 Introduction
Let $\Gamma$ be a closed smooth Riemannian manifold, ${\rm dim}\Gamma=n,$ with the metric $ds^{2}=\sum_{i,j=1}^{n}g_{ij}(u)du^{i}du^{j},$ where $u=(u^{1},\dots,u^{n})$ are local coordinates on $\Gamma$. The geodesics on $\Gamma$ are defined by the equations
$\ddot{u}^{i}+\Gamma^{i}_{jk}\dot{u}^{j}\dot{u}^{k}=0,$ (1)
where $\Gamma^{i}_{jk}$ are Christoffel symbols. In some special cases these equations can be integrated. For example, Jacobi solved the equations in the case of ellipsoids. In the general case solving these equations is a hard problem. Denote by $T^{1}\Gamma\subset T\Gamma$ the unit tangent bundle. Let $x_{0}\in\Gamma,v_{0}\in T^{1}_{x_{0}}\Gamma$ and let $\gamma(s)$ be the unique geodesic on $\Gamma$ such that $\gamma(0)=x_{0},\dot{\gamma}(0)=v_{0}$. The geodesic flow $\rho^{s}:T^{1}\Gamma\rightarrow T^{1}\Gamma$ is defined by the identity
@@
A function $F:T^{1}\Gamma\rightarrow{\mathbb{R}}$ is called a first integral if it is invariant under the flow.
The geodesic flow can also be equivalently defined by the Hamiltonian system on $T^{*}\Gamma$
@@
The geodesic flow is called Liouville–Arnold integrable if there are first integrals $G_{1}=H,G_{2},\dots,G_{n}$ functionally independent almost everywhere and $\{G_{i},G_{j}\}=0$, $i,j=1,\dots,n$; in this case, (1) can be integrated by quadratures.
There are many remarkable examples of Riemannian manifolds with integrable geodesic flows, see, e.g., [1][12] and references therein. There exist topological obstructions for integrability in the case of real analytic metrics. Kozlov’s theorem establishes that the geodesic flow on an oriented closed surface of genus $g>1$ with any real analytic metric does not admit real analytic first integrals [13]. Taimanov [14] generalized this result to the many-dimensional case. Butler [15] showed that there exist real-analytic Riemannian metrics on certain compact nilmanifolds for which the geodesic flow does not have real-analytic first integrals, while remaining integrable in the smooth category. Bolsinov and Taimanov constructed an example of real-analytic metric on a $3$-dimensional manifold whose geodesic flows are smoothly integrable and yet have positive topological entropy [16]. In general, the problem of the existence of metrics on an arbitrary manifold with integrable geodesic flows is very interesting and difficult. For example, it is unclear whether there exist metrics on closed surfaces of genus $g>1$ with Liouville–Arnold integrable geodesic flow in the smooth category. It is a very interesting question whether there exist metrics on the two-dimensional torus whose geodesic flows admit irreducible polynomial first integrals in $p_{1},p_{2}$ of degree greater than two (see, e.g., [17]).
In this paper we study the geodesic flow on the cone over an arbitrary closed $C^{3}$ Riemannian manifold $\Gamma$. By Nash embedding theorem, $\Gamma$ can be isometrically embedded in $\mathcal{P}=\{x^{N+1}=1\}\subset\mathbb{R}^{N+1}$ for some $N>n$. Let us consider the cone over $\Gamma$
$K=\{\,tp\mid p\in\Gamma,\;t\in\mathbb{R}\,\}\subset\mathbb{R}^{N+1}.$
We denote its upper and lower parts by
@@
Let $O$ be the singular point of $K$ ($t=0$). To define $TK$ it is natural to set
$T_{O}K:=K.$ (2)
We show below that if a geodesic on $K$ contains $O$ or has $O$ as a limit point, then the geodesic is a cone generatrix, i.e. a line passing through $O$ (see Lemma 2). Hence the geodesic flow
@@
Figure 1: The sphere as a caustic of the billiard inside a cone.
In [18, 19] , the billiard dynamics inside $K^{+}$ was studied in the case where $\dim\Gamma=N-1$. It was shown that the line containing any segment of a billiard trajectory in $K^{+}$ is tangent to a fixed sphere centered at $O$ (see Fig. 1). Hence the radius $r$ of this sphere serves as a first integral of the billiard system. It turns out that an analogous property holds for geodesics on a cone. For any non-generatrix geodesic on $K$, all its tangent lines are tangent to a common sphere (see Fig. 2).
Figure 2: All tangent lines of a given non-generatrix geodesic are tangent to a common sphere.
@@
Moreover, for billiards, if $\Gamma$ is $C^{3}$-smooth and strictly convex with an everywhere nondegenerate second fundamental form, the system admits a set of first integrals whose values uniquely determine every trajectory. We extend this idea to the geodesic flow on $K$, constructing a set of first integrals that uniquely determine all non-radial geodesics. Let us define
$\mathcal{T}=\left\{(x,v)\in T^{1}K\mid x\neq kv\text{ for any }k\in\mathbb{R}\right\}\subset T^{1}K.$
$\mathcal{T}$ is an open dense subspace of $T^{1}K$. Any element $(x,v)\in\mathcal{T}$ defines a non-generatrix geodesic $\gamma(s)$ with $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$. The main result of this paper is the following.
@@
There exist continuous first integrals $I^{1},\ldots,I^{2N+2}:T^{1}K\to\mathbb{R}$ of the geodesic flow on $T^{1}K$, which are $C^{1}$-smooth on $\mathcal{T}$, such that the image of the map
@@
has the property that each point in the image, except the origin in $\mathbb{R}^{2N+2}$, determines a unique geodesic on $K$, while the origin corresponds to all radial geodesics (the cone generatrices) with on $K$.
Although it is unclear when the geodesic flow on an arbitrary Riemannian manifold $\Gamma$ is integrable, on the cone over $\Gamma$ the geodesic flow is integrable in the sense that it admits enough first integrals to uniquely define almost all geodesics except for cone generatrices.
In the next theorem we state several properties of geodesics on $K$. It turns out that there is a connection between geodesics on $K$ and geodesics on
$\Sigma=K\cap\mathbb{S}^{N},$
where $\mathbb{S}^{N}\subset\mathbb{R}^{N+1}$ is the unit sphere centered at $O$. Let $\gamma(s)\subset K$ be a non-radial geodesic. Then $I>0$ and $s\in(-\infty,+\infty)$ (see Lemma 2 below). One observes that $a_{1}\gamma(a_{2}s)$, $a_{1},a_{2}\in\mathbb{R}\setminus\{0\}$, is also a geodesic. Thus, without loss of generality, we may assume that $I=1$ and $\|\gamma^{\prime}(s)\|=1$. We have:
Theorem 2.
@@
    $\tilde{\gamma}(\tilde{s})=\frac{\gamma(s)}{\|\gamma(s)\|},$ (4)
@@
The correspondence in Theorem 2 shows that non-radial geodesics on $K$ are completely determined, up to a radial scaling, by geodesic segments of fixed length on $\Sigma$, which is the underlying reason why the geodesic flow on $K$ admits the set of first integrals stated in Theorem 1.
In section 2 we will prove Lemma 1 and Theorem 2. In section 3 we will prove Theorem 1.
@@ 2 Geodesics on Cones
In this section we prove Theorem 2. We start by proving two lemmas. Lemma 1 (see Introduction) provides an essential observation about the geodesic flow on a cone; its proof is straightforward.
@@
$2\langle\gamma(s),\gamma^{\prime}(s)\rangle-2\langle\gamma(s),\gamma^{\prime}(s)\rangle\left(\langle\gamma^{\prime}(s),\gamma^{\prime}(s)\rangle+\langle\gamma(s),\gamma^{\prime\prime}(s)\rangle\right)\frac{1}{\|\gamma^{\prime}(s)\|^{2}}=0.$
@@
By Lemma 1, $I\geq 0$. The next lemma classifies geodesics on the cone according to whether $I=0$ or $I>0$.
@@
Every geodesic $\gamma(s)$ on $K$ with unit speed $\|\gamma^{\prime}(s)\|=1$, after an appropriate shift of parameter, belongs to one of the following two classes:
@@
    Radial geodesics (generatrices). These are straight lines of the form
@@
    passing through the vertex $O$. They satisfy $I=0$.
@@
    Non-radial geodesics. They satisfy $I>0$ and can be represented as
@@
    where $q(s)$ is some curve on $\Sigma$.
Proof.
@@
Since $\gamma(s)$ is a geodesic on $K$, it satisfies $\langle\gamma^{\prime\prime}(s),\gamma(s)\rangle=0$ and $\|\gamma^{\prime}(s)\|=1$. Therefore,
@@
$\gamma^{\prime}(s)=\tilde{\gamma}^{\prime}(\arctan s)\frac{1}{\sqrt{s^{2}+1}}+\tilde{\gamma}(\arctan s)\frac{s}{\sqrt{s^{2}+1}},$
@@
Thus $\gamma^{\prime\prime}(s)$ is orthogonal to $T_{\gamma(s)}K$, and therefore $\gamma(s)$ is a geodesic on $K$.
@@
Let us compute using $\|\gamma^{\prime}(s)\|=1$ and (10)
@@
Theorem 2 is proved. ∎
3 Construction of first integrals, proof of Theorem 1
In this section, we prove Theorem 1 by first constructing smooth integrals on $\mathcal{T}$ (see Lemma 3) and then extending them to continuous integrals on $T^{1}K$ using the integral $I$ (see Lemma 4).
For $(x,v)\in\mathcal{T}$, let $\gamma_{x,v}(s)$ denote the geodesic with $\gamma_{x,v}(0)=x$ and $\gamma^{\prime}_{x,v}(0)=v$. By part 1) of Theorem 2, any non-radial geodesic $\gamma_{x,v}(s)$ on $K$ touches the sphere
@@ 3 Construction of first integrals, proof of Theorem 1
at a unique point $\gamma_{x,v}(s_{0})$, where $s_{0}=s_{0}(x,v)$ is a function depends on $(x,v)$. See Fig. 3. Since the cone $K$ is embedded in $\mathbb{R}^{N+1}$, both the position vector $\gamma_{x,v}(s_{0})$ and the velocity vector $\gamma_{x,v}^{\prime}(s_{0})$ can be naturally regarded as vectors in $\mathbb{R}^{N+1}$. We may write
$\gamma_{x,v}(s_{0})=(J^{1}_{0},\dots,J^{N+1}_{0}),\qquad\gamma_{x,v}^{\prime}(s_{0})=(J^{N+2}_{0},\dots,J^{2N+2}_{0}).$
Figure 3: The geodesic $\gamma_{x,v}(s)$ touches $\mathbb{S}^{N}(\sqrt{I})$ at the unique point $\gamma_{x,v}(s_{0})$.
@@
$J^{j}(x,v):=\gamma_{x,v}^{j}(s_{0}(x,v)),\qquad j=1,\dots,N+1,$ (17)
$J^{N+1+j}(x,v):=\gamma_{x,v}^{\prime\,j}(s_{0}(x,v)),\qquad j=1,\dots,N+1.$ (18)
Here the superscript $j$ denotes the $j$-th Euclidean coordinate of a vector in $\mathbb{R}^{N+1}$. By construction, the functions $J^{1},\dots,J^{2N+2}$ are constant along each geodesic trajectory in $\mathcal{T}$ and therefore form first integrals of the system.
The next lemma establishes the smoothness of integrals $J^{1},\dots,J^{2N+2}$ on $\mathcal{T}$ and shows that they uniquely determine non-radial geodesics.
@@
The functions $J^{1},\dots,J^{2N+2}$ defined above are $C^{1}$-smooth first integrals of the geodesic flow on $\mathcal{T}$. Moreover, the image point
$\mathcal{J}=(J^{1},\dots,J^{2N+2})$
@@
Proof of Lemma 3.
The geodesic equations on $TK^{+}\cup TK^{-}$ can be written as a first-order system on the tangent bundle, which corresponds to a $C^{1}$ vector field $X$. By Lemma 2, the open subset $\mathcal{T}\subset T^{1}K^{+}\cup T^{1}K^{-}$ is invariant under the associated flow $\rho^{s}$. By the theory of ordinary differential equations, the flow $\rho^{s}(x,v)$ depends $C^{1}$-smoothly on the initial data $(x,v)\in\mathcal{T}$. Moreover, since $\partial_{s}\rho^{s}(x,v)=X(\rho^{s}(x,v))$ and the vector field $X$ is $C^{1}$, it follows that $\rho^{s}(x,v)$ is $C^{2}$ with respect to the time parameter $s$.
@@
From the proof of part 1) of Theorem 2, for each $(x,v)$, $s_{0}=s_{0}(x,v)$ is the unique time such that
@@
Since $\partial_{s}\phi(x,v,s)|_{s=s_{0}}=\|\gamma^{\prime}_{x,v}(s_{0})\|^{2}=\|v\|^{2}>0$ and $\phi$ is $C^{1}$ in $(x,v,s)$, the implicit function theorem implies that $s_{0}(x,v)$, founded from (19), is $C^{1}$ with respect to $(x,v)$.
@@
By the construction (17) and (18), it follows that $(J^{1}(x,v),\dots,J^{2N+2}(x,v))$ depend $C^{1}$-smoothly of $(x,v)$, as they are compositions of two $C^{1}$ maps. This proves the $C^{1}$-smoothness of $J^{1},\dots,J^{2N+2}$ on $\mathcal{T}$.
@@
Finally, to see that $\mathcal{J}=(J^{1},\dots,J^{2N+2})$ uniquely determines a non-radial geodesic, note that $J$ gives the position $x_{0}=(J^{1},\dots,J^{N+1})$ and velocity $v_{0}=(J^{N+2},\dots,J^{2N+2})$ at the unique parameter value where $\langle\gamma_{x,v}(s),\gamma^{\prime}_{x,v}(s)\rangle=0$. Since the geodesic initial value problem has a unique solution, this data determines a unique geodesic. ∎
Let us define

### v2

# Integrable Geodesic Flows on Cones over Riemannian Manifolds

@@
Integrable Geodesic Flows on Cones over Riemannian Manifolds
Andrey E. Mironov and Siyao Yin
Аннотация
In this paper we study the behavior of geodesics on cones over arbitrary $C^{3}$-smooth closed Riemannian manifolds. We show that the geodesic flow on such cones admits first integrals whose values uniquely determine almost all geodesics except for radial geodesics; thus, the geodesic flow is superintegrable. Moreover, we prove that the geodesic flow restricted to the open dense subset of the cotangent bundle corresponding to all non-radial trajectories is Liouville–Arnold integrable. This investigation is inspired by our recent results on Birkhoff billiards inside cones over convex manifolds where similar results hold true.
@@ 1 Introduction
Let $\Gamma$ be a closed smooth Riemannian manifold, $\operatorname{dim}\Gamma=n$, with the metric $ds^{2}=\sum_{i,j=1}^{n}g_{ij}(u)du^{i}du^{j},$ where $u=(u^{1},\dots,u^{n})$ are local coordinates on $\Gamma$. The geodesics on $\Gamma$ are defined by the equations $\ddot{u}^{i}+\Gamma^{i}_{jk}\dot{u}^{j}\dot{u}^{k}=0$, where $\Gamma^{i}_{jk}$ are Christoffel symbols. In some special cases these equations can be integrated. For example, Jacobi solved the equations in the case of ellipsoids. In the general case solving these equations is a hard problem. Denote by $T^{1}\Gamma\subset T\Gamma$ the unit tangent bundle. Let $x_{0}\in\Gamma,v_{0}\in T^{1}_{x_{0}}\Gamma$ and let $\gamma(s)$ be the unique geodesic on $\Gamma$ such that $\gamma(0)=x_{0},\dot{\gamma}(0)=v_{0}$. The geodesic flow $\rho^{s}:T^{1}\Gamma\rightarrow T^{1}\Gamma$ is defined by the identity
@@
A function $F:T^{1}\Gamma\rightarrow{\mathbb{R}}$ is called a first integral if it is invariant under the flow. The geodesic flow can also be equivalently defined by the Hamiltonian system on $T^{*}\Gamma$
@@
The geodesic flow is called Liouville–Arnold integrable if there are first integrals $G_{1}=H,G_{2},\dots,G_{n}$ functionally independent almost everywhere and $\{G_{i},G_{j}\}=0$, $i,j=1,\dots,n$; in this case, the geodesic equations can be integrated by quadratures.
There are many remarkable examples of Riemannian manifolds with integrable geodesic flows, see, e.g., [1][11] and references therein. There exist topological obstructions for integrability in the case of real analytic metrics. Kozlov’s theorem establishes that the geodesic flow on an oriented closed surface of genus $g>1$ with any real analytic metric does not admit real analytic first integrals [12]. Taimanov [13] generalized this result to the many-dimensional case. Butler [14] showed that there exist real-analytic Riemannian metrics on certain compact nilmanifolds for which the geodesic flow does not have real-analytic first integrals, while remaining integrable in the smooth category. Bolsinov and Taimanov constructed an example of real-analytic metric on a $3$-dimensional manifold whose geodesic flows are smoothly integrable and yet have positive topological entropy [15]. In general, the problem of the existence of metrics on an arbitrary manifold with integrable geodesic flows is very interesting and difficult. For example, it is unclear whether there exist metrics on closed surfaces of genus $g>1$ with Liouville–Arnold integrable geodesic flow in the smooth category. It is a very interesting question whether there exist metrics on the two-dimensional torus whose geodesic flows admit non-reducible polynomial first integrals in $p_{1},p_{2}$ of degree greater than two (see, e.g., [16]).
In this paper we study the geodesic flow on the cone over an arbitrary closed $C^{3}$ Riemannian manifold $\Gamma$, $\dim\Gamma=n$. By Nash embedding theorem, $\Gamma$ can be isometrically embedded in $\mathcal{P}=\{x^{N+1}=1\}\subset\mathbb{R}^{N+1}$ for some $N>n$. Let us consider the cone over $\Gamma$
$K=\{\,tp\mid p\in\Gamma,\;t\in\mathbb{R}\,\}\subset\mathbb{R}^{N+1},$
equipped with the Riemannian metric induced by the ambient Euclidean metric of $\mathbb{R}^{N+1}$. We denote its upper and lower parts by
@@
Let $O$ be the singular point of $K$ ($t=0$). To define $TK$ it is natural to set $T_{O}K:=K$. We show below that if a geodesic on $K$ contains $O$ or has $O$ as a limit point, then the geodesic is a cone generatrix, i.e. a line passing through $O$ (see Lemma 3). Hence the geodesic flow
@@
Рис. 1: The sphere as a caustic of the billiard inside a cone.
In [17, 18] the billiard dynamics inside $K^{+}$ was studied in the case where $\dim\Gamma=N-1$. It was shown that the line containing any segment of a billiard trajectory in $K^{+}$ is tangent to a fixed sphere centered at $O$ (see Fig. 1). Hence the radius $r$ of this sphere serves as a first integral of the billiard system. It turns out that an analogous property holds for geodesics on a cone. For any non-generatrix geodesic on $K$, all its tangent lines are tangent to a common sphere (see Fig. 2).
Рис. 2: All tangent lines of a given non-generatrix geodesic are tangent to a common sphere.
@@
Moreover, for billiards, if $\Gamma$ is $C^{3}$-smooth and strictly convex with an everywhere nondegenerate second fundamental form, the system admits a set of first integrals whose values uniquely determine every trajectory. We extend this idea to the geodesic flow on $K$, constructing a set of first integrals that uniquely determine all non-radial geodesics. The first main result of the paper is the following.
@@
There exist continuous first integrals $I^{1},\ldots,I^{2N+2}:T^{1}K\to\mathbb{R}$ of the geodesic flow on $T^{1}K$, which are $C^{1}$-smooth on $T^{1}K^{+}\cup T^{1}K^{-}$, such that the image of the map
@@
has the property that each point in the image, except the origin in $\mathbb{R}^{2N+2}$, determines a unique geodesic on $K$, while the origin corresponds to all radial geodesics on $K$.
Although it is unclear when the geodesic flow on an arbitrary Riemannian manifold $\Gamma$ is integrable, on the cone over $\Gamma$ the geodesic flow is superintegrable in the sense that it admits enough first integrals to uniquely define almost all geodesics except for cone generatrices.
In the next lemma, we state several properties of geodesics on $K$. Let
$\Sigma=K\cap\mathbb{S}^{N},\qquad\dim\Sigma=n,$
where $\mathbb{S}^{N}\subset\mathbb{R}^{N+1}$ is the unit sphere centered at $O$ and let $g_{\scriptscriptstyle\Sigma}$ denote the induced metric on $\Sigma$. The metric on the cone $K$ has the form
$g=dt^{2}+t^{2}g_{\scriptscriptstyle\Sigma},$
which is an example of a warped product metric, studied for example in [19]. In particular, Proposition 7.38 together with Remark 7.39 in [19] imply that the projection of any geodesic in $K$ onto $\Sigma$ is a pregeodesic on $\Sigma$ (that is, a geodesic up to reparametrization). In the following lemma we give an explicit formulation of this correspondence in the case of the cone and analyze the asymptotic behavior of the geodesic flow as $s\to\pm\infty$.
Suppose $\gamma(s)\subset K$ is a non-radial geodesic. Then $I>0$ and $s\in(-\infty,+\infty)$ (see Lemma 3 below). One observes that $a_{1}\gamma(a_{2}s)$, $a_{1},a_{2}\in\mathbb{R}\setminus\{0\}$, is also a geodesic. Thus, without loss of generality, in the next lemma we assume that $I=1$ and $\|\gamma^{\prime}(s)\|=1$. We have:
Lemma 2.
@@
    $\tilde{\gamma}(\tilde{s})=\frac{\gamma(s)}{\|\gamma(s)\|}\subset\Sigma,$ (2)
@@
The correspondence in Lemma 2 shows that non-radial geodesics on $K$ are completely determined, up to a radial scaling, by geodesic segments of fixed length on $\Sigma$, which is the underlying reason why the geodesic flow on $K$ admits the set of first integrals stated in Theorem 1.
In the next theorem, we will prove that the geodesic flow is Liouville–Arnold integrable when restricted to the open dense subset $\mathcal{M}\subset T^{*}K^{+}$ consisting of all cotangent vectors associated with non-radial geodesics. To describe $\mathcal{M}$, let us introduce local coordinates $(t,z)$ on $K^{+}$, where $t$ is the radial coordinate and $z$ represents coordinates on $\Sigma$. Let $(p_{0},p)$ be the coordinates on the cotangent fibers induced by the local coordinates $(t,z)$. Note that radial geodesics are characterized by $z$ being constant (i.e., the tangent vector is purely radial), which implies that the dual coordinates $p$ vanish. Therefore, the condition for a cotangent vector to correspond to a non-radial geodesic is the non-vanishing of $p$; hence the set $\mathcal{M}$ is characterized by
$\mathcal{M}=\{(t,z;p_{0},p)\in T^{*}K^{+}\mid p\neq 0\}.$
The restriction of the geodesic flow to $\mathcal{M}$ is smooth and complete (see Lemma 3). Let $H$ be the Hamiltonian generating the geodesic flow. The second main result of the paper is the following.
Theorem 2.
The geodesic flow restricted to the open dense subset $\mathcal{M}\subset T^{*}K^{+}$ is Liouville–Arnold integrable; that is, there exist $n+1$ $C^{1}$-smooth first integrals $\mathcal{F}_{0}=H,\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ on $\mathcal{M}$ that are pairwise in involution, i.e., $\{\mathcal{F}_{i},\mathcal{F}_{j}\}=0$, and functionally independent almost everywhere.
Remark 3.
If $\Gamma$ is a $C^{r}$ Riemannian manifold, then the first integrals $\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ are of class $C^{r-2}$.
By our construction, the first integrals $\mathcal{F}_{1},\dots,\mathcal{F}_{n}$ can not be defined on the subset of $T^{*}K^{+}$ that corresponds to the radial geodesics.
In section 2 we will prove Lemma 1 and Lemma 2. In section 3 we will prove Theorem 1. In section 4 we will prove Theorem 2.
@@ 2 Geodesics on Cones
We start by proving Lemma 1 (see Introduction). Lemma 1 provides an essential observation of the geodesic flow on a cone; its proof is straightforward.
@@
$2\langle\gamma(s),\gamma^{\prime}(s)\rangle-2\langle\gamma(s),\gamma^{\prime}(s)\rangle\Big(\langle\gamma^{\prime}(s),\gamma^{\prime}(s)\rangle+\langle\gamma(s),\gamma^{\prime\prime}(s)\rangle\Big)\frac{1}{\|\gamma^{\prime}(s)\|^{2}}=0.$
@@
By Lemma 1, $I\geq 0$. For our convenience before proving Lemma 2 we first prove the following lemma, which classifies geodesics on the cone according to whether $I=0$ or $I>0$.
@@
Every geodesic $\gamma(s)$ on $K$ with $\|\gamma^{\prime}(s)\|=1$, after an appropriate shift of parameter, belongs to one of the following two classes:
@@
    Radial geodesics (generatrices):
@@
    which are straight lines passing through the vertex $O$. They satisfy $I=0$.
@@
    Non-radial geodesics:
@@
    where $q(s)$ is some curve on $\Sigma$. They satisfy $I>0$.
Доказательство.
@@
Since $\gamma(s)$ is a geodesic on $K$, it satisfies $\langle\gamma^{\prime\prime}(s),\gamma(s)\rangle=0$, and by our assumption $\|\gamma^{\prime}(s)\|=1$. Therefore,
@@
$\gamma^{\prime}(s)=\tilde{\gamma}^{\prime}(\arctan s)\frac{1}{\sqrt{s^{2}+1}}+\tilde{\gamma}(\arctan s)\frac{s}{\sqrt{s^{2}+1}},$ (13)
@@
Thus $\gamma^{\prime\prime}(s)$ is orthogonal to $T_{\gamma(s)}K$, and therefore $\gamma(s)$ is a geodesic on $K$. From (13) it follows that
$\|\gamma^{\prime}(s)\|=\|\gamma^{\prime}(0)\|=\|\tilde{\gamma}^{\prime}(0)\|=1.$
@@
Using $\|\gamma^{\prime}(s)\|=1$ and (8) let us compute
@@
Lemma 2 is proved. ∎
3 Proof of Theorem 1
Let us define two subsets of $T^{1}K$:
$\mathcal{T}$ $=\left\{(x,v)\in T^{1}K\mid x\neq kv\text{ for any }k\in\mathbb{R}\right\},$
Then $\mathcal{T}\cup\mathcal{S}=T^{1}K$ and $\mathcal{T}$ is an open dense subspace of $T^{1}K$. By Lemma 3, any element $(x,v)\in\mathcal{T}$ defines a non-radial geodesic $\gamma(s)$ with $\gamma(0)=x$ and $\gamma^{\prime}(0)=v$.
In this section, we prove Theorem 1 by first constructing smooth integrals on $\mathcal{T}$ (see Lemma 4) and then modify them to obtain integrals smooth on $T^{1}K^{+}\cup T^{1}K^{-}$ and continuous on $T^{1}K$ (see Lemma 5).
3.1 Construction of first integrals
For $(x,v)\in\mathcal{T}$, let $\gamma_{x,v}(s)$ denote the geodesic with $\gamma_{x,v}(0)=x$ and $\gamma^{\prime}_{x,v}(0)=v$. From Lemma 2, it follows that any non-radial geodesic $\gamma_{x,v}(s)$ on $K$ touches the sphere
@@ 3.1 Construction of first integrals
at a unique point $\gamma_{x,v}(s_{0})$, where $s_{0}=s_{0}(x,v)$ is a function depends on $(x,v)$ (see Fig. 3). Since the cone $K$ is embedded in $\mathbb{R}^{N+1}$, both the position vector $\gamma_{x,v}(s_{0})$ and the velocity vector $\gamma_{x,v}^{\prime}(s_{0})$ can be naturally regarded as vectors in $\mathbb{R}^{N+1}$.
Рис. 3: The geodesic $\gamma_{x,v}(s)$ touches $\mathbb{S}^{N}(\sqrt{I})$ at the unique point $\gamma_{x,v}(s_{0})$.
@@
$J_{j}(x,v):=\gamma_{x,v}^{j}(s_{0}(x,v)),\qquad j=1,\dots,N+1,$ (16)
$J_{N+1+j}(x,v):=\gamma_{x,v}^{\prime\,j}(s_{0}(x,v)),\qquad j=1,\dots,N+1.$ (17)
Here the superscript $j$ denotes the $j$-th Euclidean coordinate of a vector in $\mathbb{R}^{N+1}$. By construction, the functions $J_{1},\dots,J_{2N+2}$ are constant along each geodesic trajectory in $\mathcal{T}$ and therefore form first integrals of the geodesic flow restricted to $\mathcal{T}$.
The next lemma establishes the smoothness of integrals $J_{1},\dots,J_{2N+2}$ on $\mathcal{T}$ and shows that they uniquely determine non-radial geodesics.
@@
The functions $J_{1},\dots,J_{2N+2}$ are $C^{1}$-smooth first integrals of the geodesic flow restricted to $\mathcal{T}$. Moreover, the image point
$\mathcal{J}=(J_{1},\dots,J_{2N+2})\in\mathbb{R}^{2N+2}$
@@
Доказательство.
By Lemma 3, the open subset $\mathcal{T}\subset T^{1}K^{+}\cup T^{1}K^{-}$ is invariant under the geodesic flow $\rho^{s}$. The geodesic flow restricted to $\mathcal{T}$ corresponds to a $C^{1}$ vector field on $T\mathcal{T}$:
$X(\rho^{s}(x,v))=\partial_{s}\rho^{s}(x,v).$
By the theory of ordinary differential equations, the flow $\rho^{s}(x,v)$ depends $C^{1}$-smoothly on the initial data $(x,v)\in\mathcal{T}$. Moreover, since the vector field $X$ is $C^{1}$, $\rho^{s}(x,v)$ is $C^{1}$ with respect to the parameter $s$.
@@
From the proof of part 1) of Lemma 2, for each $(x,v)$, $s_{0}=s_{0}(x,v)$ is the unique value of the parameter such that
@@
Since $\partial_{s}\phi(x,v,s)|_{s=s_{0}}=\|\gamma^{\prime}_{x,v}(s_{0})\|^{2}=\|v\|^{2}=1>0$ and $\phi$ is $C^{1}$ in $(x,v,s)$, the implicit function theorem implies that $s_{0}(x,v)$, founded from (18), is $C^{1}$ with respect to $(x,v)$ (later we will prove that $s_{0}(x,v)=-\langle x,v\rangle$, see (56)).
@@
By the construction (16) and (17), it follows that $(J_{1}(x,v),\dots,J_{2N+2}(x,v))$ depend $C^{1}$-smoothly of $(x,v)$, as they are compositions of two $C^{1}$ maps. This proves the $C^{1}$-smoothness of $J_{1},\dots,J_{2N+2}$ on $\mathcal{T}$.
@@
Finally, to see that $\mathcal{J}=(J_{1},\dots,J_{2N+2})$ uniquely determines a non-radial geodesic, note that $\mathcal{J}$ gives the position $x_{0}=(J_{1},\dots,J_{N+1})$ and velocity $v_{0}=(J_{N+2},\dots,J_{2N+2})$ at the unique parameter value where $\langle\gamma_{x,v}(s),\gamma^{\prime}_{x,v}(s)\rangle=0$. Since the geodesic initial value problem has a unique solution, this data determines a unique geodesic. ∎